# Star formation rate density as a function of galaxy mass at z 0.2 with MUSE and GAMA surveys

### v1

@@ Abstract
The star formation rate density (SFRD) is an important tool in galaxy evolution that allows us to identify at which cosmic time galaxies are more efficient at forming stars. For low-mass star-forming galaxies, the SFRD as a function of stellar mass can be straightforwardly related to the galaxy stellar mass function (GSMF). Given the uncertainty of the GSMF at the low-mass end, due to the challenges in observing dwarf galaxies, deriving the SFRD with respect to mass may be crucial to understand galaxy formation. Measurement of SFRD is more complete than number density in a cosmological volume because galaxies with higher SFR are easier to detect and characterize. In this work, the SFRD is derived using two different samples, one using the MUSE WIDE and MUSE Hubble Ultra Deep Field IFU spectroscopic surveys, and another using the GAMA spectroscopic survey. The first sample comprised a total of 27 star-forming galaxies at z 0.2 (H selected), whereas the second contained 7579 galaxies at z 0.06 (-band selected). The star formation rates are derived from measurements of the H emission line fluxes for the first sample, and using MagPhys SED fitting for the second one. The results show the behaviour of the SFRD to the lowest stellar masses of 105.5 , consistent with a constant slope (in log SFRD versus log stellar mass) and thus no turn-over in the GSMF. The results also suggest that the volume considered for the MUSE WIDE and HUDF sample is underdense.
@@
The star formation rate density (SFRD) is defined as the total star-formation rate (SFR) averaged over a given comoving volume of the universe observed at a given redshift (Madau et al., 1996). Understanding the evolution of the SFRD across cosmic history may unravel crucial insights into how galaxies grow and evolve, constraining cosmological models (Schaye et al., 2010). When measuring its contribution as a function of redshift, it allows us to understand at which cosmic time, star formation is more favourable in galaxies; whereas if analysed with respect to galaxy stellar mass, the SFRD may provide useful information on the role of galaxy mass in star formation Gilbank et al. (2010b).
@@
SFRD measurements over cosmic time have become a common practice in modern astrophysics in the past three decades. Numerous studies have estimated the shape of SFRD, with literature agreeing that it peaks at around 2-3, declines in more recent epochs, with another sharper drop at 8 (Bouwens et al., 2015). Lilly et al. (1996) pioneered the analysis of SFRD, conducting the first systematic work to combine a large and deep spectroscopic redshift survey with multi-wavelength photometry, and deriving the luminosity function and SFRD to redshifts of . Hopkins & Beacom (2006) showed how the cosmic star formation history may be disclosed by measurements of the SFRD, as it shows in which epochs star formation is more favourable. HST data was used by Bouwens et al. (2015) to push the limits of SFRD to the highest redshifts, as they derived the SFRD from UV luminosity functions to redshifts of 10, showing a sharp decrease in star formation at redshifts of 8. Madau & Dickinson (2014) reviewed over 200 SFRD measurements from the literature, reporting methodologies and results, alongside connecting the observational results with cosmological theories.
@@
Although there has been a vast number of studies on how SFRD evolves over time, the same cannot be said on how SFRD relates to other intrinsic galaxy properties. Understanding the relation between SFRD and other physical properties may give us additional information on how galaxies evolve and assemble (Madau & Dickinson, 2014). For example, James et al. (2008a) have shown that SFRD varies significantly with galaxy morphology, highlighting the importance of galaxy type in shaping star formation activity. Schulze et al. (2021) measured CCSNe rates (closely related to SFRD) as a function of galaxy mass for different types of supernovae. If analysed with respect to mass distribution, the SFRD provides useful insights of how galaxies of given masses contribute to the total star-formation-rate budget across cosmic time, revealing mass-dependent trends in star formation history (Gilbank et al., 2010a). Studies such as Drake et al. (2015) have further examined this relation, showing how star formation evolves with stellar mass between redshifts and .
@@
For low-mass star-forming galaxies, assuming that the mean specific star formation rate is approximately constant for masses smaller than , measurements of the SFRD can give an estimate of the galaxy stellar mass function (GSMF) Sedgwick et al. (2019). This assumption is confirmed by the analysis of H emission as a function of galaxy stellar mass James et al. (2008b) and by unquenched low-mass galaxies, dwarf galaxies whose cold gas reservoir has not been emptied by any physical mechanism, as they have been forming stars at a quasi-constant rate, meaning that it is reasonable to expect that their mean specific SFR is van Zee (2001).
@@
Gilbank et al. (2010b) calculated the SFRD using H, [OII], and -band luminosities from the Sloan Digital Sky Survey (SDSS) York et al. (2000) coupled with UV data from the Galaxy Evolution EXplorer (GALEX) satellite Budavári et al. (2009), deriving the SFRD with respect to stellar mass down to masses of . However, the derivation of the SFRD at lower masses has proven to be a challenge due to the extremely low surface brightness of dwarf galaxies Cross & Driver (2002). Sedgwick et al. (2019) introduced a novel approach, using core-collapse supernovae (CCSNe) as "signposts" to low surface brightness galaxies to constrain their abundance and contribution to the total star formation rate density (SFRD) at lower masses. Rates are converted to SFRDs using the expected CCSNe per unit mass of star formation. The work from Sedgwick et al. (2019) suggests that a greater fraction of star formation occurs in low-mass, low-surface brightness galaxies than previously thought. Nevertheless, more sensitive surveys are needed to fully capture the star formation activity and extend SFRD measurements to smaller and fainter systems.
@@
Previous studies of SFRD have used used emission lines and broadband photometric tracers to extend SFRD measurements to low stellar masses across cosmic time, with the H emission line being one of the best indicators of star formation in a galaxy, since young, massive stars with ages 20 Myr produce copious amounts of ionising photons that ionise the surrounding gas Davies et al. (2016). This is a great advantage when calculating the SFR of a galaxy, as only young stars are taken into account and thus it is largely independent of SF history Kennicutt Jr (1998). Additionally, compared to other emission lines, H is the most directly proportional to the ionizing UV stellar spectra at 912 Å and it is not as susceptible to the metal fraction present in the gas Tresse & Maddox (1998). A limitation of using H as a star-formation rate indicator is that the aperture based spectroscopy only probes the central regions of nearby galaxies Davies et al. (2016), but this can be corrected with the appropriate calibration or with integral-field unit (IFU) spectroscopy.
@@
As modern IFU surveys can provide complete spatial and spectral coverage, detecting emission lines even for the faintest and most diffuse sources that lack continuum, they can be used to probe the faintest galaxy populations, extending SFRD measurements to the lowest masses (Bacon et al., 2015; Herenz et al., 2017). The MUSE Hubble Ultra Deep Field (HUDF) Bacon et al. (2023) and MUSE Wide Urrutia et al. (2019) are optical IFU spectroscopic surveys from the MUSE collaboration that can reach unprecedented depths compared to similar spectroscopic surveys.
@@
In this work, the low-mass galaxy population at redshift 0.2 is characterised using MUSE Wide and MUSE HUDF, deriving their SFRs through the analysis of their emission line fluxes. Using this information, the SFRD with respect to stellar mass of the population for the combined MUSE sample is calculated. The SFRD is also derived using the Galaxy and Mass Assembly (GAMA) survey, a shallower, but much larger spectroscopic survey. The results suggest that the low-mass galaxy population contributes to the cosmic star formation rate density, albeit at lower levels than more massive galaxies. A description of the data sets used is presented in § 2, the principles of the SFRD measurements are presented in § 3, sample selection and processing in § 4, and the main results in § 5; with discussion and summary in the final sections. An flat- cosmology with is assumed for distances and volumes.
@@ 2.1 MUSE Wide
The MUSE Wide Survey is a large spectroscopic survey that uses the Multi-Unit Spectroscopic Explorer (MUSE) instrument on the Very Large Telescope (VLT) to observe galaxies in the distant universe (Bacon et al., 2010). The MUSE Wide survey is a blind, 3D spectroscopic survey that provides, like many other extragalactic surveys, a “wedding-cake” approach, and is expected to cover 100 × 1 arcmin2 MUSE fields by the end of the survey Urrutia et al. (2019). MUSE-Wide mainly covers parts of the CDFS and COSMOS regions that were previously mapped by HST in several bands to intermediate depths, by GOODS-South in the optical Giavalisco et al. (2004) and by CANDELS in the near infrared Grogin et al. (2011). The survey is designed to cover the whole field of view continuously, so that it is not restricted to a photometric pre-selection for identification and classification of objects in the sky Bacon et al. (2017).
The first data release (DR1) consists of 44 CANDELS-CDFS fields, with observations carried out in nominal mode (each spectrum spans from 4750 – 9350 Å in wavelength range), and 0.2”×0.2” spatial and 1.25 Å wavelength sampling, which is the default for MUSE Urrutia et al. (2019). The survey used a combination of spectroscopic and photometric measurements to derive the stellar masses of the galaxies in their catalogue. The photometry was taken from Skelton et al. (2014), whilst the software FAST (Fitting and Assessment of Synthetic Templates, Kriek et al. 2018) was used for the spectral energy distribution (SED) fitting model. FAST determines the best fit parameters using minimization from a set of model SEDs and an analysis grid describing several stellar population models.
The catalogues used in this study were retrieved from the MUSE Wide DR1 database, which is available on their website (MUSEWIDE_DR1). The two catalogues used in this work MUSE, one being the MUSE WIDE main table (musewide_dr1.mw_44fields_main_table), from which the columns containing galaxy properties such as redshift and stellar mass, and their associated source with the Skelton et al. (2014) catalogue, were extracted (columns names respectively: z, stellar_mass, skelton_id, skelton_sep). The other catalogue used is the MUSE WIDE emission line table (musewide_dr1.mw_44fields_emline_table), containing information about line fluxes at different Kron radius apertures (column used f_2kron, which comprises values for emission line fluxes extracted in an aperture equivalent to Kron radius 2). The emission line and source identification was performed using LSDCat (Herenz & Wisotzki, 2017), an automated detection package for emission lines in wide-field integral-field spectroscopic datacubes. The full description of the data release and catalogues is reported in the documentation from Urrutia et al. (2019) and Herenz & Wisotzki (2017).
@@ 2.2 MUSE Hubble Ultra Deep Field
From the MUSE HUDF survey, the data from the DR2 main table (dr2_main_09) is used for the analysis. The table contains galaxy properties like redshift and stellar mass derived using MagPhys, which can be accessed via their respective columns ’Z’ and ’MASS_MAG’. For each galaxy, the table also contains values for emission line fluxes, with the H flux values used being stored in the ’HALPHA_EMI_FLUX’ column. For the MUSE HUDF survey, the source detection and classification was performed using the blind detection software ORIGIN, capable of detecting faint line emitters in MUSE datacubes (Mary et al., 2020). The dataset is publicly available and may be retrieved from the collaboration website (MUSEHUDF_DR2). For full description of the survey and catalogues, please refer to Bacon et al. (2023).
@@ 2.3 GAMA
The Galaxy And Mass Assembly (GAMA) survey (Liske et al., 2015; Baldry et al., 2018; Driver et al., 2022) is a spectroscopic redshift and multi-wavelength photometric survey designed to study galaxy evolution. The survey is divided into five regions: three equatorial regions (G09, G12, G15) each of 60.0 deg2 and main survey limit of rAB 19.8 mag, and two southern regions of 55.7 deg2 (G02) and 50.6 deg2 (G23). Here, we use the equatorial regions, which cover an area of 179 deg2 after accounting for masking around bright stars.
The fourth data release of GAMA (DR4) is the final release (Driver et al., 2022). Compared to MUSE Wide and MUSE HUDF, GAMA is shallower. Moreover, it relies on fiber spectroscopy, which may result in an overestimation of emission line fluxes and SFRs for small or low surface brightness galaxies when applying aperture corrections (Richards et al., 2016). However, GAMA benefits from a much wider area and larger catalogue, comprising 21-band photometric data with spectroscopic line emission measurements for star formation indicators OII and , containing information for stellar mass estimates of 198223 galaxies, and measured redshifts for 196402.
The survey data is organised into data management units (DMUs). The DMUs used in this work are: SpecCat, containing the spectra and redshifts from all the curated spectroscopic data of the GAMA survey, with the primary choice of redshift extracted using the automatic code AUTOZ (Baldry et al., 2014); SpecLineSFR, comprising line flux and equivalent width measurements for GAMA II spectra (Gordon et al., 2016); StellarMasses, which provides measurements of the total stellar mass of the sources (Taylor et al., 2011); MagPhys (Da Cunha et al., 2008), that contains estimates of a number of key parameters including stellar mass, dust mass, and star-formation rate (SFR) (Driver et al., 2018). For the descriptions of other DMUs, please refer to Baldry et al. (2018) and Driver et al. (2022).

### v2

# Star formation rate density as a function of galaxy mass at with MUSE and GAMA surveys

@@ Abstract
The star formation rate density (SFRD) is an important tool in galaxy evolution that allows us to identify at which cosmic time galaxies are more efficient at forming stars. For low-mass star-forming galaxies, the SFRD as a function of stellar mass can be straightforwardly related to the galaxy stellar mass function (GSMF). Given the uncertainty of the GSMF at the low-mass end, due to the challenges in observing dwarf galaxies, deriving the SFRD with respect to mass may be crucial to understand galaxy formation. Measurement of SFRD is more complete than number density in a cosmological volume because galaxies with higher SFR are easier to detect and characterize. In this work, the SFRD is derived using two different samples, one using the MUSE WIDE and MUSE Hubble Ultra Deep Field IFU spectroscopic surveys, and another using the GAMA spectroscopic survey. The first sample comprised a total of 27 star-forming galaxies at (H selected), whereas the second contained 7579 galaxies at (-band selected). The star formation rates are derived from measurements of the H emission line fluxes for the first sample, and using MagPhys SED fitting for the second one. The results show the behaviour of the SFRD to the lowest stellar masses of , consistent with a constant slope (in log SFRD versus log stellar mass) and thus no turn-over in the GSMF.
@@
The star formation rate density (SFRD) is defined as the total star-formation rate (SFR) averaged over a given comoving volume of the universe observed at a given redshift (Madau etĀ al., 1996). Understanding the evolution of the SFRD across cosmic history may unravel crucial insights into how galaxies grow and evolve, constraining cosmological models (Schaye etĀ al., 2010). When measuring its contribution as a function of redshift, it allows us to understand at which cosmic time, star formation is more favourable in galaxies; whereas if analysed with respect to galaxy stellar mass, the SFRD may provide useful information on the role of galaxy mass in star formation Gilbank etĀ al. (2010b).
@@
SFRD measurements over cosmic time have become a common practice in modern astrophysics in the past three decades. Numerous studies have estimated the shape of SFRD, with literature agreeing that it peaks at around 2-3, declines in more recent epochs, with another sharper drop at (Bouwens etĀ al., 2015). Lilly etĀ al. (1996) pioneered the analysis of SFRD, conducting the first systematic work to combine a large and deep spectroscopic redshift survey with multi-wavelength photometry, and deriving the luminosity function and SFRD to redshifts of . Hopkins & Beacom (2006) showed how the cosmic star formation history may be disclosed by measurements of the SFRD, as it shows in which epochs star formation is more favourable. HST data was used by Bouwens etĀ al. (2015) to push the limits of SFRD to the highest redshifts, as they derived the SFRD from UV luminosity functions to redshifts of , showing a sharp decrease in star formation at redshifts of . Madau & Dickinson (2014) reviewed over 200 SFRD measurements from the literature, reporting methodologies and results, alongside connecting the observational results with cosmological theories.
@@
Although there has been a vast number of studies on how SFRD evolves over time, the same cannot be said on how SFRD relates to other intrinsic galaxy properties. Understanding the relation between SFRD and other physical properties may give us additional information on how galaxies evolve and assemble (Madau & Dickinson, 2014). For example, James etĀ al. (2008a) have shown that SFRD varies significantly with galaxy morphology, highlighting the importance of galaxy type in shaping star formation activity. Schulze etĀ al. (2021) measured CCSNe rates (closely related to SFRD) as a function of galaxy mass for different types of supernovae. If analysed with respect to mass distribution, the SFRD provides useful insights of how galaxies of given masses contribute to the total star-formation-rate budget across cosmic time, revealing mass-dependent trends in star formation history (Gilbank etĀ al., 2010a; Khostovan etĀ al., 2024). Studies such as Drake etĀ al. (2015) have further examined this relation, showing how star formation evolves with stellar mass between redshifts and .
@@
For low-mass star-forming galaxies, assuming that the mean specific star formation rate is approximately constant for masses smaller than , measurements of the SFRD can give an estimate of the galaxy stellar mass function (GSMF) Sedgwick etĀ al. (2019). This assumption is confirmed by the analysis of H emission as a function of galaxy stellar mass James etĀ al. (2008b) and by unquenched low-mass galaxies, dwarf galaxies whose cold gas reservoir has not been emptied by any physical mechanism, as they have been forming stars at a quasi-constant rate, meaning that it is reasonable to expect that their mean specific SFR is van Zee (2001).
@@
Gilbank et al. (2010b) calculated the SFRD using H, [OII], and -band luminosities from the Sloan Digital Sky Survey (SDSS) York et al. (2000) coupled with UV data from the Galaxy Evolution EXplorer (GALEX) satellite BudavÔri et al. (2009), deriving the SFRD with respect to stellar mass down to masses of . However, the derivation of the SFRD at lower masses has proven to be a challenge due to the extremely low surface brightness of dwarf galaxies Cross & Driver (2002). Sedgwick et al. (2019) introduced a novel approach, using core-collapse supernovae (CCSNe) as "signposts" to low surface brightness galaxies to constrain their abundance and contribution to the total star formation rate density (SFRD) at lower masses. Rates are converted to SFRDs using the expected CCSNe per unit mass of star formation. The work from Sedgwick et al. (2019) suggests that a greater fraction of star formation occurs in low-mass, low-surface brightness galaxies than previously thought. Nevertheless, more sensitive surveys are needed to fully capture the star formation activity and extend SFRD measurements to smaller and fainter systems.
@@
Previous studies of SFRD have used used emission lines and broadband photometric tracers to extend SFRD measurements to low stellar masses across cosmic time, with the H emission line being one of the best indicators of star formation in a galaxy, since young, massive stars with ages 20 Myr produce copious amounts of ionising photons that ionise the surrounding gas Davies etĀ al. (2016). This is a great advantage when calculating the SFR of a galaxy, as only young stars are taken into account and thus it is largely independent of SF history Kennicutt (1998). Additionally, compared to other emission lines, H is the most directly proportional to the ionizing UV stellar spectra at 912 Ć Ā and it is not as susceptible to the metal fraction present in the gas Tresse & Maddox (1998). A limitation of using H as a star-formation rate indicator is that the aperture based spectroscopy only probes the central regions of nearby galaxies Davies etĀ al. (2016), but this can be corrected with the appropriate calibration or with integral-field unit (IFU) spectroscopy.
@@
As modern IFU surveys can provide complete spatial and spectral coverage, detecting emission lines even for the faintest and most diffuse sources that lack continuum, they can be used to probe the faintest galaxy populations, extending SFRD measurements to the lowest masses (Bacon etĀ al., 2015; Herenz etĀ al., 2017). The MUSE Hubble Ultra Deep Field (HUDF) Bacon etĀ al. (2023) and MUSE Wide Urrutia etĀ al. (2019) are optical IFU spectroscopic surveys from the MUSE collaboration that can reach unprecedented depths compared to similar spectroscopic surveys.
@@
In this work, the low-mass galaxy population at redshift 0.2 is characterised using MUSE Wide and MUSE HUDF, deriving their SFRs through the analysis of their H emission line fluxes. Using this information, the SFRD with respect to stellar mass of the population for the combined MUSE sample is calculated. The SFRD is also derived using the Galaxy and Mass Assembly (GAMA) survey. GAMA is a much larger spectroscopic survey but is selected using broadband SDSS photometry and is expected to be significantly incomplete to low-surface-brightness galaxies below about (Baldry etĀ al., 2012). Results of both approaches are presented to determine the contribution of the low-mass galaxy population to the cosmic star formation rate density.
A description of the data sets used is presented in § 2, the principles of the SFRD measurements are presented in § 3, sample selection and processing in § 4, and the main results in § 5; with discussion and summary in the final sections. An flat- cosmology with is assumed for distances and volumes.
@@ 2.1 MUSE Wide
The MUSE Wide Survey is a large spectroscopic survey that uses the Multi-Unit Spectroscopic Explorer (MUSE) instrument on the Very Large Telescope (VLT) to observe galaxies in the distant universe (Bacon etĀ al., 2010). The MUSE Wide survey is a blind, 3D spectroscopic survey that provides, like many other extragalactic surveys, a āwedding-cakeā approach, and is expected to cover 100 Ć 1 arcmin2 MUSE fields by the end of the survey Urrutia etĀ al. (2019). MUSE-Wide mainly covers parts of the CDFS and COSMOS regions that were previously mapped by HST in several bands to intermediate depths, by GOODS-South in the optical Giavalisco etĀ al. (2004) and by CANDELS in the near infrared (Grogin etĀ al., 2011; Koekemoer etĀ al., 2011). The survey is designed to cover the whole field of view continuously, so that it is not restricted to a photometric pre-selection for identification and classification of objects in the sky Bacon etĀ al. (2017).
The first data release (DR1) consists of 44 CANDELS-CDFS fields, with observations carried out in nominal mode (each spectrum spans from 4750 ā 9350 Ć in wavelength range), and 0.2āĆ0.2ā spatial and 1.25 Ć wavelength sampling, which is the default for MUSE Urrutia etĀ al. (2019). The survey used a combination of spectroscopic and photometric measurements to derive the stellar masses of the galaxies in their catalogue. The photometry was taken from Skelton etĀ al. (2014), whilst the software FAST (Fitting and Assessment of Synthetic Templates, Kriek etĀ al. 2018) was used for the spectral energy distribution (SED) fitting model. FAST determines the best fit parameters using minimization from a set of model SEDs and an analysis grid describing several stellar population models.
The catalogues used in this study were retrieved from the MUSE Wide DR1 database, which is available on their website (MUSEWIDE_DR1). The two catalogues used in this work MUSE, one being the MUSE WIDE main table (musewide_dr1.mw_44fields_main_table), from which the columns containing galaxy properties such as redshift and stellar mass, and their associated source with the Skelton etĀ al. (2014) catalogue, were extracted (columns names respectively: z, stellar_mass, skelton_id, skelton_sep). The other catalogue used is the MUSE WIDE emission line table (musewide_dr1.mw_44fields_emline_table), containing information about line fluxes at different Kron radius apertures (column used f_2kron, which comprises values for emission line fluxes extracted in an aperture equivalent to Kron radius 2). The emission line and source identification was performed using LSDCat (Herenz & Wisotzki, 2017), an automated detection package for emission lines in wide-field integral-field spectroscopic datacubes. The full description of the data release and catalogues is reported in the documentation from Urrutia etĀ al. (2019) and Herenz & Wisotzki (2017).
@@ 2.2 MUSE Hubble Ultra Deep Field
From the MUSE HUDF survey, the data from the DR2 main table (dr2_main_09) is used for the analysis. The table contains galaxy properties like redshift and stellar mass derived using MagPhys, which can be accessed via their respective columns āZā and āMASS_MAGā. For each galaxy, the table also contains values for emission line fluxes, with the H flux values used being stored in the āHALPHA_EMI_FLUXā column. For the MUSE HUDF survey, the source detection and classification was performed using the blind detection software ORIGIN, capable of detecting faint line emitters in MUSE datacubes (Mary etĀ al., 2020). The dataset is publicly available and may be retrieved from the collaboration website (MUSEHUDF_DR2). For full description of the survey and catalogues, please refer to Bacon etĀ al. (2023).
@@ 2.3 GAMA
The Galaxy And Mass Assembly (GAMA) survey (Liske etĀ al., 2015; Baldry etĀ al., 2018; Driver etĀ al., 2022) is a spectroscopic redshift and multi-wavelength photometric survey designed to study galaxy evolution. The survey is divided into five regions: three equatorial regions (G09, G12, G15) each of 60.0 deg2 and main survey limit of rAB 19.8 mag, and two southern regions of 55.7 deg2 (G02) and 50.6 deg2 (G23). Here, we use the equatorial regions, which cover an area of 179 deg2 after accounting for masking around bright stars.
The fourth data release of GAMA (DR4) is the final release (Driver etĀ al., 2022). Compared to MUSE Wide and MUSE HUDF, GAMA is shallower. Moreover, it relies on fiber spectroscopy, which may result in an overestimation of emission line fluxes and SFRs for small or low surface brightness galaxies when applying aperture corrections (Richards etĀ al., 2016). However, GAMA benefits from a much wider area and larger catalogue, comprising 21-band photometric data with spectroscopic line emission measurements for star formation indicators OII and H, containing information for stellar mass estimates of 198223 galaxies, and measured redshifts for 196402.
The survey data is organised into data management units (DMUs). The DMUs used in this work are: SpecCat, containing the spectra and redshifts from all the curated spectroscopic data of the GAMA survey, with the primary choice of redshift extracted using the automatic code AUTOZ (Baldry etĀ al., 2014); StellarMasses, which provides measurements of the total stellar mass of the sources (Taylor etĀ al., 2011); MagPhys (DaĀ Cunha etĀ al., 2008), that contains estimates of a number of key parameters including stellar mass, dust mass, and star-formation rate (SFR) (Driver etĀ al., 2018). For the descriptions of other DMUs, please refer to Baldry etĀ al. (2018) and Driver etĀ al. (2022). Note that we prefer the MagPhys estimates of the SFRs compared to the H fluxes (Hopkins etĀ al., 2013; Gordon etĀ al., 2016) because, in the latter case, the fluxes are indirectly derived from equivalent widths and -band photometry which can result in unverified high-SFR outliers (Davies etĀ al., 2016) that significantly impact SFRD measurements.